OpenStudy (anonymous):

Use rational exponents to write 4√2*3√7 as a single radical expression. Choose the correct expression: A.)12√19208 B.)12√14 C.)12√105413504 D.)1/19208^12

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Because: \[4\sqrt{2} \times 3\sqrt{7} = 4 \times 3 \times \sqrt{2}\sqrt{7} = 12\sqrt{}\]

OpenStudy (anonymous):

\[12\sqrt14{}\]**

OpenStudy (anonymous):

B

OpenStudy (anonymous):

THANK YOU!!